Can someone please help me out on this?
Determine the radius of a circle if a 24-inch chord is 9 inches from the center.
Thanks a lot!

Answer by rothauserc(4718)

Use pythagorean formula
9^2 + 12^2 = r^2 where r is the radius
r^2 = 225
r = 15 inches